PART 4: To answer questions 10, 11 and 12 , please read and use the following context.

Consider a monopolist that practices third-degree price discrimination (also called explicit price discrimination). For the first segment of the market, the demand for the firm’s product is given by:

P1=100-2Q1

Where the subscript 1 is used to refer to the fact this is the first segment. For the second segment, the demand for the firm’s product is given by:

P2=50-Q2

Where the subscript 2 is used to refer to the fact that this is the second segment. The cost function of the firm is:

C=20Q

Question 10: True or false, the marginal revenues from each market segment are equal to each other.

Question 11: What quantity will the firm sell to the first segment and at what price?

Question 12: What quantity will the firm sell to the second segment and at what price?
Answer #1

TO Market! Market 2 P = 100-20, P = 50-Q TR,= 100 - 202 TR = 90 Q-222 MR = 100- 4Q MR₂ = 50-222 Line, MR, ZMR2, so given stat